# HEIGHT: HEterogeneous Interaction GrapH Transformer for Robot Navigation in Crowded and Constrained Environments

### v1

@@ VII-A Sim2real through Real2sim
To overcome sim2real gaps, the design of simulation pipelines needs to be guided by the constraints of hardware and environments in the real-world. On one hand, we determine the input representation of HEIGHT based on what could be obtained from sensors and off-the-shelf perception modules and how accurate they are. We find that intermediate features, such as detected human positions and processed point clouds, reduces sim2real gaps. On the other hand, we also ensure the consistency of the simulation and real-world, such as the robot action space, whenever possible. These design choices allow our policy to generalize to different simulation environments and deploy to challenging real-world scenarios.
However, although we have minimized the sim2real gaps through real2sim, certain gaps still exist. In real-world experiments, the difficulty of the task is reduced and the agility of the robot policy is only partially transferred from simulation. To further align the simulation and the real-world, we plan to explore the following directions for future work: (1) developing a more natural pedestrian model to replace the ORCA humans in the simulator, (2) revising our pipeline to enable self-supervised RL fine-tuning in the real-world [60, 61], (3) using a small amount of real-world data to automatically optimize the parameters of our simulator to match the real-world environment [62, 63].
@@ VII-B Scene representation
A good scene representation is tailored to the needs of its downstream task. Besides the above sim2real considerations, our scene representation is split due the different nature of humans and obstacles for robot collision avoidance. The size of humans are small and their shapes are simple. Therefore, to avoid humans, the robot only needs to treat them as circles with a heading direction. In contrast, obstacles have larger and more complicated surfaces. The part of the obstacle contours that faces the robot is more important for collision avoidance. Therefore, point clouds are the most intuitive way to represent such useful information about obstacles. Our experiments in Sec. V show that this split scene representation reduces robot collision avoidance with both dynamic and static obstacles.
A side effect of our scene abstraction is the loss of detailed information such as gaits of humans and 3D shapes of obstacles. However, we argue that this is a tradeoff to minimize sim2real gaps with limited simulation tools and computation resources. Another side effect is the cascading errors between the perception modules and robot policy, such as inaccurate human detections or robot localization. To this end, it is worth exploring the joint optimization of the whole robotic stack from perception to control [64, 65].
@@ VII-C Structured neural network
Model-based approaches require low-fidelity data yet heavily rely on assumptions. In contrast, end-to-end learning approaches need few assumptions yet require high-fidelity data. Our structured learning method combines the best of both worlds: It requires low-fidelity data yet relies on minimal assumptions. By injecting structures into the neural network, we decompose a complex problem into smaller and relatively independent components. Note that our decomposition does not break the gradient flow, which keeps HEIGHT end-to-end trainable. We propose a principled way for network architecture design, increasing the transparency of these black-boxes. Our experiments demonstrate that the structured network outperforms both model-based methods and RL-based methods without structures, which empirically proves the effectiveness of structure learning for interactive tasks with multiple heterogeneous participants.
@@ VII-D Training method
Deep RL enables the robot to explore the environment and learn meaningful behaviors through trial and error. Without heuristics or demonstrations, the robot has to collect reward signals to improve its policy. Consequently, simulation development with a randomization scheme and a good reward design are indispensable to the performance of our method. Nevertheless, our method is subjective to the inherent limitations of RL, such as low training data efficiency and difficulties in long-horizon tasks. In future work, combining the RL policy with other traditional planners or imitation learning has the potential to alleviate these problems [36].
@@ VIII Conclusion
Acknowledgements
We thank Zhe Huang for helpful discussions and code of the A* planner. We thank Aamir Hasan and Rutav Shah for feedback on paper drafts. We thank members in Human-Centered Autonomy Lab who participated in real-world experiments.

### v2

@@ VII-A Sim2real through Real2sim
Lessons learned: To overcome sim2real gaps, the design of simulation pipelines needs to be guided by the constraints of hardware and environments in the real-world. On one hand, we determine the input representation of HEIGHT based on what could be obtained from sensors and off-the-shelf perception modules and how accurate they are. We find that intermediate features, such as detected human positions and processed point clouds, reduces sim2real gaps. On the other hand, we also ensure the consistency of the simulation and real-world, such as the robot action space, whenever possible. These design choices allow our policy to generalize to different simulation environments and deploy to challenging real-world scenarios.
Limitation and future work: However, although we have minimized the sim2real gaps through real2sim, certain gaps still exist. Before sim2real transfer, the policy needs to be finetined in simulations that mimic the target environments (Fig.5(d) and (e)). Despite this, in real-world experiments, the difficulty of the task is reduced and the agility of the robot policy is only partially transferred from simulation. To further align the simulation and the real-world, we plan to explore the following directions for future work: (1) developing a more natural pedestrian model to replace the ORCA humans in the simulator, (2) revising our pipeline to enable failure detections and self-supervised RL fine-tuning in the real-world [62, 63], (3) using a small amount of real-world data to automatically optimize the parameters of our simulator to match the real-world environment [64, 65].
@@ VII-B Scene representation
Lessons learned: A good scene representation should be tailored to the needs of its downstream task. Besides the above sim2real considerations, our scene representation is split due the different nature of humans and obstacles for robot collision avoidance. The size of humans are small and their shapes are simple. Therefore, to avoid humans, the robot only needs to treat them as circles with a heading direction. In contrast, obstacles have larger and more complicated surfaces. The part of the obstacle contours that faces the robot is more important for collision avoidance. Therefore, point clouds are the most intuitive way to represent such useful information about obstacles. Our experiments in Sec. I show that this split scene representation reduces robot collision avoidance with both dynamic and static obstacles.
Limitation and future work: A side effect of our scene abstraction is the loss of detailed information such as gaits of humans and 3D shapes of obstacles. However, we argue that this is a tradeoff to minimize sim2real gaps with limited simulation tools and computation resources. Another side effect is the cascading errors between the perception modules and robot policy, such as inaccurate human detections or robot localization. A future direction to automate the design of inputs and outputs is the joint optimization of the whole robotic stack from perception to control [66, 67].
@@ VII-C Structured neural network
Lessons learned: Model-based approaches require low-fidelity data yet heavily rely on assumptions. In contrast, end-to-end learning approaches need few assumptions yet require high-fidelity data. Our structured learning method combines the best of both worlds: It requires low-fidelity data yet relies on minimal assumptions. By injecting structures into the neural network, we decompose a complex problem into smaller and relatively independent components. Note that our decomposition does not break the gradient flow, which keeps HEIGHT end-to-end trainable. We propose a principled way for network architecture design, increasing the transparency of these black-boxes. Our experiments demonstrate that the structured network outperforms both model-based methods and RL-based methods without structures, which empirically proves the effectiveness of structure learning for interactive tasks with multiple heterogeneous participants.
Limitation and future work: The structured neural network utilizes the inductive bias behind human behaviors to solve specific problems. However, without re-training, the aforementioned network may not generalize to other human-robot interaction (HRI) scenarios such as autonomous driving. In other fields such as computer vision and natural language processing, foundation models that are trained on large-scale diverse data have shown promise in generalization. Inspired by the standardized benchmarks of autonomous driving and robot manipulation [68, 69], we aim to contribute to the benchmarking efforts of human-centered tasks.
@@ VII-D Training method
Lessons learned: Deep RL enables the robot to explore the environment and learn meaningful behaviors through trial and error. Without heuristics or demonstrations, the robot has to collect reward signals to improve its policy. Consequently, simulation development with a randomization scheme and a good reward design are indispensable to the performance of our method.
Limitation and future work: Nevertheless, our method is subjective to the inherent limitations of RL, such as the need to build simulations, low training data efficiency, and difficulties in long-horizon tasks. In future work, combining the RL policy with other traditional planners or imitation learning has the potential to alleviate these problems [70].
@@ VIII Conclusion
Appendix
Figure 10: Visualizations of HH and RH attention weights of HEIGHT. The color of lines indicate the weight values (see the colorbar on the right). For simplicity, we only include the line if the weight value is greater than 0.1. We observe that the interactions that likely to affect robot navigation in the near future have high attention values. Examples include humans that are approaching the robot and humans that block the robot’s path to the goal.
-A Simulation experiments – Baselines
  • $\textrm{A}^{*}$+ CNN [7] is a hybrid method. With a map of the environment, $\textrm{A}^{*}$ is the global planner and generates 6 waypoints in the beginning of an episode. The inputs to the robot RL policy are a 2D LiDAR point cloud, the robot state, the waypoints, and the robot’s final goal. No human detections are used and human features are included in the point cloud. The point cloud is passed through a CNN local planner, whose output is concatenated with the embedding of other inputs. In addition to the reward function in Eq. 1, the robot receives a heading direction award if it moves towards the goal while being collision-free.
  • DRL-VO [30] is a hybrid method with a pure pursuit algorithm as the global planner and a RL policy as the local planner. The inputs to the planners are the positions and velocities of humans encoded in an occupancy map, a 2D LiDAR point cloud, and the robot state. The inputs are passed into a CNN network without explicit interaction modeling.
-B Simulation experiments – Training
-C Simulation experiments – Additional insights
Effectiveness of RL planning: In Table II, DWA is a model-based approach that determines robot actions based on only the current state, which leads to the worst performance. Thus, reactive policies are not sufficient to solve our problem, justifying the necessity of long-sighted planning. In contrast, $\textrm{A}^{*}$+ CNN combines planning and RL, yet relies on occupancy maps with only obstacles and not humans. As humans move, the increasingly inaccurate occupancy map reduces the optimality of waypoints, which negatively affects overall navigation. On the other hand, RL learns to maximize the expected long-term returns based on both the current and historical states of all components in the scene. Consequently, the remaining 3 RL methods, especially HomoGAT and HEIGHT, outperform $\textrm{A}^{*}$+ CNN in most metrics and in most environments. Thus, to ensure the best navigation performance, it is important to optimize the entire planning system based on the task setting instead of optimizing only a part of it.
Difficulty of scenarios: From Table II and Table III, we observe that, besides the overfitted DS-RNN, all methods show the same trend of performance change in the 5 environments: Less crowded $>$ Less constrained $>$ More constrained $>$ More crowded. Obviously, adding more humans or obstacles increases task difficulty. But interestingly, the change in the number of humans has a larger effect on the task difficulty than the change in the number of obstacles. This phenomenon shows that avoiding collisions with humans is more difficult than avoiding obstacles in nature.
-D Real-world experiment setup
The distance between the starting and the goal position of the robot ranges from 6m to 11m. The pedestrians were told to react naturally to the robot based on their own preferences. In some testing episodes, other pedestrians who were unaware of our experiment also engaged with the robot.